OpenStudy (itsmichelle29):

A house cost $120,000 when it was purchased. The value of the house increases by 10% each year. Find the rate of growth each month and select the correct answer below. 0.797% 0.00797% 0.833% 0.00833%

OpenStudy (itsmichelle29):

I medal and fan .... I think its C

OpenStudy (princeharryyy):

correct :)

OpenStudy (itsmichelle29):

CAN YOU HELP ME WITH ONE MORE

OpenStudy (princeharryyy):

yes!

OpenStudy (itsmichelle29):

Ashley solved the exponential equation 3x+1 = 15 and her work is shown below. What is the first step she did incorrectly? Step 1: log 3^(x+1) = log15 Step 2: (x + 1)log 3 = log15 Step 3: x + 1 = log 15/log 3 Step 4: x+1 = 2.708050/1.098612 Step 5: x + 1 = 1.418858 Step 6: x = 0.418858

OpenStudy (itsmichelle29):

thank you .... IDK if its step 3 or step 4

OpenStudy (princeharryyy):

Step 4: x+1 = 2.708050/1.098612

OpenStudy (princeharryyy):

values of log are wrong!

OpenStudy (itsmichelle29):

what about 3

OpenStudy (itsmichelle29):

thanks

OpenStudy (princeharryyy):

three is even.

OpenStudy (princeharryyy):

it's valid to write it like that.